# A Systematic Survey on Large Language Models for Algorithm Design

### v1

@@ 4.2.4. Reinforcement Learning
Reinforcement learning learns to make decisions by interacting with an environment to maximize cumulative rewards. Recent works have explored innovative applications and enhancements of this method. For instance, GPTswarm (Zhuge et al., 2024) and the work by Duan et al. (2023) leverage LLMs and reinforcement learning techniques to optimize code, reducing complexity and enhancing efficiency. Similarly, Liu et al. (2024f) present a novel LLM-guided search framework (LLM-GS) aimed at addressing the sample inefficiency in state-of-the-art programmatic reinforcement learning methods by utilizing the programming expertise of LLMs to boost search efficiency.
Additionally, LLMs have been utilized in the design of reinforcement learning systems, particularly in the creation of reward-shaping functions. Bhambri et al. (2024) tackle the issue of sample inefficiency by proposing a framework called MEDIC, which uses LLMs to generate a guide policy for constructing reward-shaping functions. This framework shows substantial improvements in sample complexity for RL agents across various domains in the BabyAI environment.
@@ 4.3. Population-based Search
Evolutionary search is the main tool investigated due to its effectiveness and robustness in complex search space (O’Reilly and Hemberg, 2024; Wu et al., 2024a). According to the number of objectives, it can be roughly categorized into single-objective evolutionary search and multi-objective evolutionary search.
@@ 4.3.1. Single-objective Evolutionary Search
Evolutionary algorithms are the major solutions when integrating LLM for algorithm design (Liu et al., 2024d; Romera-Paredes et al., 2024). Through maintaining a population of diverse individuals, they usually search the space effectively.
The study by Yang et al. (2023) represents an initial exploration of utilizing LLMs as optimizers for algorithm design. It suggests leveraging the in-context learning capability of LLMs to generate new solutions for target problems based on existing evaluated solutions. Although it does not explicitly mention evolutionary search, the search paradigm behind it can be regarded as a population-based evolutionary search. Liu et al. (2023a) introduce LLMs as evolutionary operators for addressing multi-objective problems by decomposing them into single-objective sub-problems and employing LLMs as black-box search operators for each sub-problem. This decomposition-based approach enables LLMs to tackle multi-objective problems indirectly, overcoming their inherent challenges in comprehending such problems. Additionally, Liu et al. (2024b) explore the use of LLMs as evolutionary operators in conventional EA, guiding them in solution selection, crossover, and mutation processes. Moreover, Brahmachary et al. (2024) propose a population-based evolutionary framework comprising exploration and exploitation pools, with LLMs facilitating solution generation for both pools and enabling communication between them during optimization. Lange et al. (2024) investigate the application of LLMs in designing evolution strategies, introducing a prompting strategy to enhance mean statistic performance.
In contrast to the integration of LLM to generate hard-coded solutions, the effective evolutionary search of prompt and code generation poses challenges due to the lack of suitable diversity measurement, selection, and population management methods on the complex search space. The majority of workers adopt a greedy way (Liu et al., 2024d; Chen et al., 2024a; van Stein and Bäck, 2024; Li et al., 2024b; Narin, 2024; Guo et al., 2024c; Yao et al., 2024a). In these works, a population of individuals is maintained and only the ones with better fitness will survive. Romera-Paredes et al. (2024) adopt a multi-island evolutionary algorithm to explore a diverse population with a dynamically adjusted population size. Moreover, Wong et al. (2024a) explore the potential of leveraging generative artificial intelligence models to create digital artifacts for creative and engineering applications, utilizing an multi-factorial evolutionary algorithm to drive an LLM. The proposed LLM2FEA approach demonstrates the ability to generate novel and practical designs, as evidenced by experimental results in the context of 3D aerodynamic design.
To enable effective diversity control over the population, Lehman et al. (2023) and Hu et al. (2024) have investigated quality diversity methods (Pugh et al., 2016). Specifically, they utilize the MAP-Elites algorithm with a grid of niches to generate diverse and high-quality solutions. A pre-existing solution is evaluated and placed into the map, with subsequent iterations perturbing solutions within inhabited niches. New candidate solutions are evaluated and assigned niches based on behavior characterization, becoming champions if they outperform current inhabitants or fill unfilled niches. This process gradually fills the map with improved solutions over time.
@@ 4.3.2. Multi-objective Evolutionary Search
Liu et al. (2023a) propose MOEA/D-LMO to use LLM to solve continuous multi-objective optimization problems. Benefiting from the decomposition-based framework, the in-context learning of LLM is used to generate new solutions for each subproblem. For multi-objective LLM instruction generation, Yang and Li (2023) develop InstOptima, which utilizes an LLM to simulate instruction operators and improve instruction quality to simultaneously optimize performance, length, and perplexity of the instruction. Moreover, Morris et al. (2024) introduce guided evolution for the development of neural architectures. It adopts NSGA-II and optimizes both accuracy and model size. MEoH (Yao et al., 2024b) introduce a dominance-dissimilarity score for multi-objective heuristic design. In contrast to conventional MOEAs, it achieves a good balance of exploration and exploitation in the heuristic search space.
@@ 4.4. Uncertainty-guided Search
Uncertainty-guided search harnesses the advantages of BOED and BO frameworks within the search process incorporating LLMs. It first establishes a prior that describes the initial belief about the parameters of interest and then utilizes uncertainty-driven strategies over the derived posterior to iteratively update the belief, ultimately achieving efficient decision-making during the search process in a sample-efficient manner.
This search paradigm with uncertainty has demonstrated its superiority in some scenarios. For example, in tasks of eliciting human preferences information using LLMs (Li et al., 2023b; Piriyakulkij et al., 2023), Handa et al. (2024) introduce OPEN that takes advantage of LLMs and Bayesian Optimal Experiment Design (BOED). Specifically, LLMs are utilized to extract environment-relevant features and translate abstract feature queries into natural language questions to interact with users through conversations, BOED is employed to identify the most informative query for efficiently learning preference information. Experimental studies have demonstrated the superiority of the OPEN approach in content recommendation tasks when compared to both LLM- and BOED-based preference elicitation methods. To improve the multi-turn and decision-theoretic reasoning capability over a set of arbitrary items, Austin et al. (2024) formulate the natural language-based preference elicitation within a Bayesian optimization framework. The proposed PEBOL algorithm maintains a probabilistic prior over user preferences in the form of Beta distributions. The generation of queries is facilitated by LLM-based acquisition functions incorporating strategies like Thompson Sampling and Upper Confidence Bound, which aim to learn user preferences efficiently balancing exploration and exploitation. In the domain of LLM decoding, Grosse et al. (2024) propose a probabilistic model named ULTS. This method frames the decoding process as a tree-search problem and identifies an optimal path in the LLM search tree by sequential decision-making under uncertainty. It firstly establishes a prior belief on the softmax outputs of the LLM and then utilizes the implied posterior samples to guide the search process in a non-myopic fashion.
Among these search methods mentioned in this section, evolutionary algorithms (EAs) are the most used search methods in LLM-based algorithm design (Cai et al., 2024), Table 1 summarizes and compares different EA-based methods in terms of their algorithmic components, i.e., prompt strategy, search strategy, and population management. For prompt strategy, we identify One-Shot prompting (OS), Few-Shot prompting (FS) (Brown et al., 2020), Chain-of-Thought (CoT) (Wei et al., 2022), and Reflection (RF) (Brown et al., 2020). For search methods, we identify Genetic Algorithm (GA), Differential Evolution (DE), Evolutionary Strategy (ES), Multi-objective Evolutionary Algorithm (MOEA) and others. The results reveal that: 1) FS and CoT are the most commonly used prompt engineering strategies. 2) The majority of works use a simple genetic algorithm with greedy population management. 3) LLMaO and LLMaD are more commonly investigated than the other two categories.
Table 1. Existing LLM-based Methods with Evolutionary Computation for Algorithm Design
Approaches Method Prompt Strategy Search Strategy Population Management
LLMaP LAEA (Hao et al., 2024) FS, CoT GA Greedy
LLMaO EvoLLM (Lange et al., 2024) FS ES All
EvolPrompt (Guo et al., 2023) FS, CoT GA, DE Greedy, Fixed Size
EvolPrompting (Chen et al., 2024a) FS, CoT GA Greedy
Instoptima (Yang and Li, 2023) FS, CoT MOEA Greedy
MOEA/D-LMO (Liu et al., 2023a) FS MOEA MOEA/D, Fixed Size
LMEA (Liu et al., 2023a) CoT GA Greedy
LEO (Brahmachary et al., 2024) FS, CoT Two-pop GA Greedy
GE (Morris et al., 2024) FS, CoT MOEA Greedy
LLMaD ELM (Lehman et al., 2023) FS MAP-Elites MAP-Elites
AEL (Liu et al., 2023b) FS, CoT GA Greedy
Eureka (Ma et al., 2024b) FS, RF GA (1+1) Greedy
FunSearch (Romera-Paredes et al., 2024) FS Multi-island GA Greedy, Dynamic Size
EoH (Liu et al., 2024d) FS, OS, CoT GA Greedy, Fixed Size
L-AutoDA (Guo et al., 2024c) FS, OS, CoT GA Greedy
EvolCAF (Yao et al., 2024a) FS, OS, CoT GA Greedy
ReEvo (Ye et al., 2024a) CoT, RF GA Greedy, Fixed Size
1+1 (Zhang et al., 2024b) OS GA (1+1) Greedy, Fixed Size
LLaMEA (van Stein and Bäck, 2024) CoT GA Greedy
Auto MC-Reward (Li et al., 2024b) CoT GA (1+1) Greedy
EROM (Narin, 2024) FS, RF GA (1+1) Greedy
@@ 5. Prompt Strategies
Fig. 9(a) depicts the percentage of domain or pre-trained LLMs used in the literature. Among them, over 80% chose to use the pre-trained model without any specific finetuning, and about 10% fine-tuned the pre-trained model on domain datasets (Cummins et al., 2024) in which 4.4% are trained from scratch. Fig. 9(c) illustrates the top-used LLMs. LLM4AD papers show a strong preference for GPT models. GPT-4 and GPT-3.5 are the most-used LLMs, which in total take around 50%. Llama-2 are the most used open-source LLM. The development of an effective domain LLMs poses challenges to most algorithm designers. We will provide more discussions in the future work section.
Once we have the pre-trained LLMs, prompt engineering is significant for effectively integrating LLMs in algorithm design. We introduce the main prompt engineering methods including zero-shot, few-shot, chain-of-thought, self-consistency, and reflection (Sahoo et al., 2024) used in LLM4AD papers. Fig. 9(b) shows the number of papers involving different prompt engineering techniques. Few-shot, zero-shot, and chain-of-thought are the most commonly used ones. Many works involve more than one prompt engineering technique.
@@ 5.1. Zero-Shot
Zero-Shot (ZS) learning enables a model to comprehend and perform tasks without prior specific training on those tasks. In algorithm design, zero-shot prompting allows direct requests to the LLM for solutions or responses, examples including conversational replies (Sekulić et al., 2024), assertion design (Miftah et al., 2024), reinforcement learning guidance (Zhang et al., 2024e), and molecule generation (Ye, 2024). The model leverages its pre-trained knowledge to generate these responses. Although this method is advantageous for swiftly producing solutions based on general knowledge, it may not consistently deliver tailored responses to the nuanced demands of complex algorithmic challenges (Zhang et al., 2024b).
@@ 5.2. Few-Shot
Few-Shot (FS) learning involves providing the model with a handful of examples to illustrate the task before it attempts to solve a similar problem. This approach enhances the model’s understanding of the context and its ability to tailor responses to specific tasks. In algorithm design, this method proves particularly effective by presenting the model with examples of algorithms (Liu et al., 2023b), solutions (Yang et al., 2024), prompts (Wang et al., 2023b; Tang et al., 2024c), and codes (Ma et al., 2024a; Morris et al., 2024). The examples used in the prompt may be manually-designed seeds (Romera-Paredes et al., 2024; Hu et al., 2024) or derived from LLM-generated designs (Liu et al., 2024d; Wong et al., 2024a; Ma et al., 2024c; Morris et al., 2024; Wang et al., 2023b; Tang et al., 2024c; Custode et al., 2024). These samples are typically sorted to suggest a preference to enhance performance (Yang et al., 2024; Romera-Paredes et al., 2024).
@@ 5.5. Reflection
Reflection in prompt engineering involves asking the model to critique or evaluate its own responses or solutions. After generating an algorithm, the model can be prompted to reflect on the efficiency, potential flaws, or improvements. This not only helps in refining the algorithm but also in understanding deeper nuances, such as space-time trade-offs, which are crucial for advanced algorithm design. Ma et al. (2024d) investigate both implicit and explicit reflection in LLM-driven prompt optimization, which lets LLM analyze the errors and generate a reflection or feedback regarding the current prompt. An additional study by Ye et al. (2024b) incorporates short-term and long-term reflections in heuristic design, while Zhang et al. (2024c) focus on self-evaluation and reasoning over results from network detection. Furthermore, both Ma et al. (2024b) and Narin (2024) adopt LLM for RL reward function design with a reward reflection that monitors the scalar values of all reward components and the task fitness function at various policy checkpoints during training.
@@ 6.1. Optimization
In this subsection, we focus on discussing the practical applications of LLM in algorithm design for optimization. We categorize the existing literature into combinatorial optimization, continuous optimization, Bayesian optimization, prompt optimization, and optimization modeling based on the specific types of optimization tasks. We then proceed to compare the various roles played by LLMs, the prompt strategies utilized, and the specific problems or tasks to which they are applied. This comparative analysis is summarized in Table 2, where we list the names of the frameworks or methods proposed by the authors. For studies that do not explicitly name their methods, we assign appropriate designations in our article and denote them with asterisks for easy reference (e.g., MH-LLM* for (Sartori et al., 2024)).
@@ 6.1.1. Combinatorial Optimization
In the domain of Combinatorial Optimization (CO), the design of effective algorithm heuristics has been a significant area of interest for a long time. The emergence of LLMs has now provided new avenues for automatically generating CO algorithms for diverse application tasks.
The Traveling Salesman Problem (TSP) stands out as one of the most renowned combinatorial optimization challenges, involving the quest for the shortest route to visit all specified locations exactly once and return to the starting point. Given its NP-hard nature, heuristic algorithms are commonly employed to tackle this problem. Some recent work leverages LLMs to evolve algorithms within Evolutionary Computation (EC) framework, such as AEL (Liu et al., 2023b), ReEvo (Ye et al., 2024a) and EoH (Liu et al., 2024d). Differently, OPRO (Yang et al., 2024) employs LLMs as optimizers with a proposed meta-prompt, in which the solution-score pairs with task descriptions are added in each optimization step. Additionally, LMEA (Liu et al., 2024c) investigates the utilization of LLMs as evolutionary combinatorial optimizers for generating offspring solutions, wherein a self-adaptation mechanism is introduced to balance exploration and exploitation. The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing constraints related to vehicle capacity. To address this challenge, MLLM (Huang et al., 2024c) devises a multi-modal LLM-based framework with textual and visual inputs to enhance optimization performance.
In addition to routing problems, the Bin Packing Problem (BPP) is another well-studied problem with diverse applications. The objective of the BPP is to pack items of varying sizes into the smallest number of fixed-sized bins. FunSearch (Romera-Paredes et al., 2024) utilizes an evolutionary approach to evolve heuristics for the BPP. The study showcases that FunSearch is capable of uncovering enhanced heuristics when compared to traditional strategies such as first fit and best fit in addressing the challenges posed by the BPP. In comparison to FunSearch, EoH (Liu et al., 2024d) designs heuristics that not only achieve competitive performance but also demonstrate superior generalization across various problem sizes and capacities, requiring significantly fewer queries. Furthermore, in addition to the previously discussed research, numerous important combinatorial optimization problems have been extensively studied to validate the effectiveness of their proposed methods. Examples include the cap set problem in FunSearch (Romera-Paredes et al., 2024), Flow Job Shop Scheduling Problem (FSSP) in EoH (Liu et al., 2024d), and social networks problem in MH-LLM* (Sartori et al., 2024).

### v2

@@ 4.2.4. Reinforcement Learning
Reinforcement learning learns to make decisions by interacting with an environment to maximize rewards. GPTswarm (Zhuge et al., 2024) and Duan et al. (2023) leverage LLMs and RL techniques to optimize code, reducing complexity and enhancing efficiency. Similarly, Liu et al. (2024f) present a novel LLM-guided search framework (LLM-GS) aimed at addressing the sample inefficiency in state-of-the-art programmatic RL methods by utilizing the programming expertise of LLMs to boost search efficiency. Additionally, LLMs have been utilized in the design of RL systems, particularly in the creation of reward-shaping functions. Moreover, Bhambri et al. (2024) propose a framework called MEDIC to tackle the issue of sample inefficiency. They use LLMs to generate a guide policy for constructing reward-shaping functions.
@@ 4.3. Population-based Search
Population-based evolutionary search is the main tool investigated in LLM4AD papers due to its effectiveness and robustness in complex search space (O’Reilly and Hemberg, 2024; Wu et al., 2024a). The majority of these works use a simple genetic algorithm with greedy population management Liu et al. (2024d). Some of them investigate advanced population management including multi-island Romera-Paredes et al. (2024), quality-diversity Hu et al. (2024), and dominance-dissimilarity measurement Yao et al. (2024b). According to the number of objectives, they can be roughly categorized into single-objective search and multi-objective search.
@@ 4.3.1. Single-objective Evolutionary Search
Liu et al. (2024c) explore the use of LLMs as evolutionary operators in conventional EA, guiding them in solution selection, crossover, and mutation processes. Moreover, Brahmachary et al. (2024) propose a population-based evolutionary framework comprising exploration and exploitation pools, with LLMs facilitating solution generation for both pools and enabling communication between them during optimization. Lange et al. (2024) investigate the application of LLMs in designing evolution strategies, introducing a prompting strategy to enhance mean statistic performance.
On algorithm design tasks involving text and code search, the complicated search space poses challenges for diversity control and population management. The majority of workers adopt a greedy way (Liu et al., 2024d; Chen et al., 2024a; van Stein and Bäck, 2024; Li et al., 2024b; Narin, 2024; Guo et al., 2024b; Yao et al., 2024a). In these works, a population of individuals is maintained and only the ones with better fitness will survive. Romera-Paredes et al. (2024) adopt a multi-island evolutionary algorithm to explore a diverse population with a dynamically adjusted population size. Moreover, Wong et al. (2024a) use LLMs to create digital artifacts for creative and engineering applications, utilizing a multi-factorial evolutionary algorithm to drive an LLM. Quality diversity methods (Pugh et al., 2016) have also been investigated for this purpose, Lehman et al. (2023) and Hu et al. (2024) utilize the MAP-Elites algorithm with a grid of niches to generate diverse and high-quality solutions.
@@ 4.3.2. Multi-objective Evolutionary Search
Liu et al. (2023a) propose MOEA/D-LMO to use LLM to solve continuous multi-objective optimization problems. Benefiting from the decomposition-based framework, the in-context learning of LLM is used to generate new solutions for each subproblem. For multi-objective LLM instruction generation, Yang and Li (2023) develop InstOptima, which utilizes an LLM to simulate instruction operators and improve instruction quality to simultaneously optimize performance, length, and perplexity of the instruction. Moreover, Morris et al. (2024) introduce guided evolution for neural architecture design. It adopts NSGA-II and optimizes both accuracy and model size. In order to enhance the diversity control in multiobjective heuristic design, MEoH (Yao et al., 2024b) introduce a dominance-dissimilarity score for multi-objective heuristic design, achieving a good balance of exploration and exploitation in the heuristic search space.
@@ 4.4. Uncertainty-guided Search
Uncertainty-guided search harnesses the strengths of Bayesian Optimal Experiment Design (BOED) or Bayesian Optimization (BO) frameworks within the search process incorporating LLMs. It first establishes a prior that describes the initial belief about the parameters of interest and then employs uncertainty-driven strategies over the derived posterior to iteratively update the belief, ultimately enabling efficient decision-making during the search process in a sample-efficient manner.
This search paradigm with uncertainty has demonstrated its superiority in some scenarios. For example, Handa et al. (2024) introduce OPEN using LLMs to extract environment-relevant features and translate abstract feature queries into natural language questions to interact with users through conversations. Meanwhile, BOED is employed to identify the most informative query for efficiently learning preference information. To improve the multi-turn and decision-theoretic reasoning capabilities over a set of arbitrary items, Austin et al. (2024) formulate the natural language-based preference elicitation within a BO framework. The proposed PEBOL algorithm establishes a probabilistic prior over user preferences in the form of Beta distributions. In the domain of LLM decoding, Grosse et al. (2024) propose a probabilistic model named ULTS. This method frames the decoding process as a tree-search problem and aims to find the optimal path in the LLM search tree by sequential decision-making under uncertainty.
@@ 5. Prompt Strategies
Fig. 8(a) depicts the percentage of domain or pre-trained LLMs used in the literature. Among them, over 80% chose to use the pre-trained model without any specific finetuning, and about 10% fine-tuned the pre-trained model on domain datasets (Cummins et al., 2024) in which 4.4% are trained from scratch. Fig. 8(c) illustrates the top-used LLMs. LLM4AD papers show a strong preference for GPT models. GPT-4 and GPT-3.5 are the most-used LLMs, which in total take around 50%. Llama-2 is the most used open-source LLM. Once we have the pre-trained LLMs, prompt engineering is significant for effectively integrating LLMs in algorithm design. We introduce the main prompt engineering methods including zero-shot, few-shot, chain-of-thought, self-consistency, and reflection (Sahoo et al., 2024) used in LLM4AD papers. Fig. 8(b) shows the number of papers involving different prompt engineering techniques.
@@ 5.1. Zero-Shot
Zero-Shot (ZS) prompting enables a model to comprehend and perform tasks without prior specific training on those tasks. In algorithm design, zero-shot prompting allows direct requests to the LLM for solutions or responses, examples including conversational replies (Sekulić et al., 2024), assertion design (Miftah et al., 2024), RL guidance (Zhang et al., 2024e), and molecule generation (Ye, 2024). The model leverages its pre-trained knowledge to generate these responses. Although this method is advantageous for swiftly producing solutions based on general knowledge, it may not consistently deliver tailored responses to the nuanced demands of complex algorithmic challenges (Zhang et al., 2024b).
@@ 5.2. Few-Shot
Few-Shot (FS) prompting involves providing the model with a handful of examples to illustrate the task before it attempts to solve a similar problem. This approach enhances the model’s understanding of the context and its ability to tailor responses to specific tasks. In algorithm design, this method proves particularly effective by presenting the model with examples of algorithms (Liu et al., 2023b), solutions (Yang et al., 2024), prompts (Wang et al., 2023b; Tang et al., 2024c), and codes (Ma et al., 2024a; Morris et al., 2024). The examples used in the prompt may be manually-designed seeds (Romera-Paredes et al., 2024; Hu et al., 2024) or derived from LLM-generated designs (Liu et al., 2024d; Wong et al., 2024a; Ma et al., 2024c; Morris et al., 2024; Wang et al., 2023b; Tang et al., 2024c; Custode et al., 2024). These samples are typically sorted to suggest a preference to enhance performance (Yang et al., 2024; Romera-Paredes et al., 2024).
@@ 5.5. Reflection
Reflection in prompt engineering involves asking the model to critique or evaluate its own responses or solutions. After generating an algorithm/solution, the model can be prompted to reflect on the efficiency, potential flaws, or improvements. Ma et al. (2024d) investigate both implicit and explicit reflection in LLM-driven prompt optimization, which lets LLM analyze the errors and generate a reflection or feedback regarding the current prompt. An additional study by Ye et al. (2024b) incorporates short-term and long-term reflections in heuristic design, while Zhang et al. (2024c) focus on self-evaluation and reasoning over results from network detection. Furthermore, both Ma et al. (2024b) and Narin (2024) adopt LLM for RL reward function design with a reward reflection that monitors the scalar values of all reward components and the task fitness function at various policy checkpoints during training.
@@ 6.1. Optimization
In this subsection, we delve into the practical applications of LLMs in algorithm design for optimization. We categorize the existing literature into combinatorial optimization, continuous optimization, Bayesian optimization, prompt optimization, and optimization modeling based on the specific domains of optimization applications. Then we proceed to compare the various roles played by LLMs, the prompt strategies employed, and the specific problems or tasks to which they are applied. The comparative analysis is summarized in Table 1, where we list the names of the frameworks or methods proposed by the authors. For studies that do not explicitly name their methods, we assign appropriate designations in our article and denote them with asterisks for easy reference (e.g., MH-LLM* for (Sartori et al., 2024)).
@@ 6.1.1. Combinatorial Optimization
In the domain of Combinatorial Optimization (CO), automatic algorithm heuristics design has been a significant area of interest for a long time. The Traveling Salesman Problem (TSP) stands out as one of the most renowned CO problems, involving the quest for the shortest route to visit all specified locations exactly once and return to the starting point. Some recent work leverages LLMs to evolve algorithms within Evolutionary Computation (EC) framework, such as AEL (Liu et al., 2023b), ReEvo (Ye et al., 2024a) and EoH (Liu et al., 2024d). Differently, OPRO (Yang et al., 2024) employs LLMs as optimizers with a proposed meta-prompt, in which the solution-score pairs with task descriptions are added in each optimization step. Additionally, LMEA (Liu et al., 2024b) investigates the utilization of LLMs as evolutionary combinatorial optimizers for generating offspring solutions, wherein a self-adaptation mechanism is introduced to balance exploration and exploitation. The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing constraints related to vehicle capacity. To address this challenge, MLLM (Huang et al., 2024c) devises a multi-modal LLM-based framework with textual and visual inputs to enhance optimization performance.
In addition to routing problems, other combinatorial optimization problems that have been investigated include cap set (Romera-Paredes et al., 2024), bin packing (Liu et al., 2024d), flow job shop scheduling (Liu et al., 2024d) and social networks problems (Sartori et al., 2024).